OpenStudy (anonymous):

If 8 is added to twice a number and this sum is multiplied by 3, the result is the same as if the number is multiplied by 4 and 10 is added to the product. What is the number?

OpenStudy (anonymous):

3(8 + 2x) = 4x + 10 24 + 6x = 4x + 10 24 + 2x = 10 2x = -14 x = -7.

OpenStudy (anonymous):

thank you!